OpenStudy (anonymous):

y = ln(5x^3 – 2e^2x + 1)^4 find the derivative dont know how to set up the rule for this derivative

OpenStudy (anonymous):

thats too long O.O \[y = \ln(junk) \Rightarrow y' = \frac{1}{junk}*(junk)'\]

myininaya (myininaya):

\[y=\ln[g(x)]^4=4\ln[g(x)] => y'=4*\frac{g'(x)}{g(x)}\]

OpenStudy (anonymous):

g(x)....junk....same stuff lol

myininaya (myininaya):

lol

myininaya (myininaya):

g(x)=junk^4

OpenStudy (anonymous):

joe the first part of your setup is all needed?

myininaya (myininaya):

wait g(x) didn't include 4

myininaya (myininaya):

lol

OpenStudy (anonymous):

myininaya's formula is good, i would use hers. to turn that annoying power into something thats not very annoying at all.

myininaya (myininaya):

but joe's formula is good because its all you really need to know to do these

myininaya (myininaya):

joe's is more general than mine

OpenStudy (anonymous):

now i am confused but thank you both

OpenStudy (anonymous):

lol >.< sry sry

myininaya (myininaya):

lol they are basically the same i just have a contant times (junk)'/junk

myininaya (myininaya):

where my junk almost equal's joe's junk

myininaya (myininaya):

lol ok i will stop

OpenStudy (anonymous):

hahahaha

OpenStudy (anonymous):

i wish i could give more medals lol

myininaya (myininaya):

yep